OpenStudy (anonymous):

chris is responsible for arranging chairs at the meeting. there are 72 chairs and he wants to have twice as many chairs in each row as he has in each column how many chairs should he put in each row? how many rows does he need?

OpenStudy (anonymous):

r=2c r+c=72 this is what you need to solve r=24 c=48 I guess I mixed up the r and c :)

OpenStudy (anonymous):

and it is not even correct :D r=2c r*c=72 c^2=36 c=6 r=12